OpenStudy (anonymous):

Find the integral of t/(sqrt(2t^2-1)) dt. (Answer: (1/2)*sqrt(2t^2-1)+C)

OpenStudy (btaylor):

I believe you need to do some substitution. I'll try my best to put it in the equation form: \[\int\limits \frac{t}{\sqrt{2t^2-1}}dt\] Let \(u=2t^2-1\). Then, \(du=4t~dt\). If we solve for \(dt\), we get \(dt=\frac{1}{4t} du\). Substituting for \(dt\), we get: \[\int\limits \frac{t}{\sqrt{u}}\frac{1}{4t}du\] Can you solve it from here?

OpenStudy (anonymous):

Thank you.

OpenStudy (btaylor):

Anytime.